OpenStudy (anonymous):

Let f(x)=1/square root of x and g(x)=1-x^2. Find f(g(x)).

OpenStudy (jdoe0001):

\(\bf f(x) = \cfrac{1}{\sqrt{x}}\\ g(x) = \color{red}{1-x^2}\\ f(g(x)) = \cfrac{1}{\sqrt{(\color{red}{1-x^2})}}\)

OpenStudy (anonymous):

Tenk u!

OpenStudy (jdoe0001):

yw